# Prediction intervals for overdispersed binomial endpoints and their application to toxicological historical control data

### v3

@@ 5.2 Coverage probabilities for the simulation setting inspired by long term carcinogenicity studies
The heuristical methods (Fig. 5) behaved in a similar fashion than in the MNT-setting: With rising amount of historical control groups, the future observation is always covered (Fig. 5A). In the absence of overdispersion, the np-Chart (Fig. 5B) yields coverage probabilities close to the nominal level, but with a rising amount of overdispersion, its coverage probabilities decrease below 0.6.
With a rising amount of historical control groups (at least 20), the coverage probabilities of the HCL computed by the mean $\pm$ 2 SD (Fig. 5C) approach the nominal level, if the underlying distribution is relatively symmetrical (low overdispersion and relatively high proportion). With increasing right-skewness of the underlying distribution (increasing overdispersion and decreasing proportion) and /or a low amount of historical control groups ($H<20$) the HCL computed by the mean $\pm$ 2 SD yield coverage probabilities below the nominal level that can decrease to approximately 85 %.
@@ 6 Application of control limits
Table 3: Control limits (CL) for the mortality of male B6C3F1-mice obtained in negative controls of long term carcinogenicity studies at the National Toxicology Program.
@@ 7 Discussion
The validation of a concurrent control group based on prediction intervals is recommended by several authors (Menssen 2023, Kluxen et al. 2021) and since shortly, seems to be preferred by the European Food Safety Agency (EFSA 2024) over other (heuristical) methods. Nevertheless, since literature for the calculation of prediction intervals for clustered dichotomous data is relatively scarce, the practical application of HCL tends to rely on several heuristics, such as the application of the mean $\pm$ 2 SD (Rotolo et al. 2021, Prato et al. 2023, de Kort et al. 2020, Kluxen et al. 2021) or limits from Shewhart control charts (Lovell et al. 2018, Dertinger et al. 2023). Our simulations have shown that heuristical methods do not control the statistical error to a satisfactory level and hence, can not be recommended for the practical validation of a current control group.
Contrary to this, we could show, that prediction intervals satisfactorily control the statistical error, if calculated based on a model that match the statistical properties of the HCD. Several prediction intervals are available from the literature that are based on the assumption of independent binomial distributed observations (see e.g. Meeker et al. 2017 or Wang 2008), but ignore the clustered structure of the HCD (and hence the possibility for between-study overdispersion). In the context of HCD from long term carcinogenicity studies, Menssen and Schaarschmidt 2019 showed, that these intervals fail to approach the nominal coverage probability in the presence of overdispersion. This demonstrates that between-study overdispersion is a crucial part of the underlying data generating process and hence needs to be considered for HCL calculation.
The magnitude of overdispersion reflects the between-study variation which is commonly present in toxicological HCD and hence, can be found in several published HC data sets in the literature (Tarone 1982a, Tarone 1982b, Levy et al. 2019, Menssen et al. 2024, Carlus et a.l 2013, Dertinger et al. 2023, Tug et al. 2024, see supplementary materials section 4). However, in practice, systematic between-study variation might be induced by a mixture of both, controllable causes such as different handling of experimental units by different lab personnel, as well as uncontrollable causes such as the genetic variation inherent in the population of experimental units. This means that in practice, overdispersed HCD should be handled with care in a way that the magnitude of overdispersion reflects mainly the biological variability that is inherent of the biological process under observation, rather than the variability that is induced by controllable causes.
Both of the calibrated prediction intervals are improved versions of the quasi-binomial prediction interval of Menssen and Schaarschmidt 2019 which rely on a bootstrap procedure that adapts the formerly symmetrical prediction interval to possible skewness of the underlying distribution. Note, that this bootstrap calibration procedure has already been applied successfully for the calculation of prediction intervals drawn from random effects models (Menssen and Schaarschmidt 2022) or for overdispersed count data (Menssen et al. 2024). This demonstrates that, as long as one is able to bootstrap new observations from a model and the prediction standard error $\sqrt{\widehat{var}(\hat{y}^{*})+\widehat{var}(Y^{*})}$ can be formulated, the presented bootstrap calibration procedure provides a standardized framework for the calculation of prediction intervals for various distributional assumptions and/or models. An implementation of the bisection used in step 8 and 9 of the bootstrap calibration is provided via the function bisection() of the R package predint.
@@
Contrary to the Bayesian hierarchical models, the Bayesian GLMM did not have convergence problems in the LTC-setting. But, the applied GLMM mainly yield prediction intervals with coverage probabilities above the nominal level (except for a higher amount of overdispersion). This is, because such models a priori assume overdispersion or, equivalently, additional non-zero between-study variability to be present, and hence provide estimates for the respective parameters, even if between-study variability is absent in the the data-generating process.
An alternative to a fully Bayesian approach might be the application of an empirical Bayes procedure in which the hyperparameters domain is pre-estimated from the HCD, similar to the approaches of Tarone 1982a who incorporated HCD in a trend test or Kitsche et al. 2012 who exploit HCD to be accounted for in a Dunnett-type mean comparison. Anyhow, to the authors knowledge, an empirical Bayes procedure for the calculation of prediction intervals for overdispersed binomial data does not exist and might be a subject for future research.
@@
It is noteworthy, that the coverage probability of the calibrated intervals is mainly effected by the amount of available historical information, defined by the numbers of experimental units within each control group $n_{h}$ and number of historical studies $H$. In most of the cases the amount of overdispersion or the magnitude of the binomial proportion is of minor importance (as long as one accepts that the lower border can become zero for low proportions and / or high overdispersion such that the desired 95 % prediction interval practically results in a 97.5 % upper prediction limit).
This means, that for a high cluster size (as in the MNT-setting) both calibrated prediction intervals yield coverage probabilities satisfactorily close to the nominal level even for only five historical control groups. Contrary, for smaller cluster sizes (as in the LTC-setting) the number of historical control groups has to be slightly higher (say 10) to enable coverage probabilities close enough to the nominal level. Furthermore, note that with decreasing cluster size ($n_{h}$ and $n^{*}$), the discreteness of the underlying distribution rises and hence, the nominal 0.95 coverage probability can not completely be achieved in all of the possible cases (Agresti 1998).
@@
Since the bootstrap calibration procedure presented above has already been applied to enable the calculation of prediction intervals drawn from linear random effects models (Menssen and Schaarschmdit 2022) or in order to predict overdispersed count data (Menssen et al. 2024), it can be seen as a general and flexible tool for the calculation of prediction intervals. Hence, for the future it is planned to exploit its potential to enable the calculation for a broader range of models / distributions. Furthermore it is possible to extend the algorithm to enable the calculation of simultaneous prediction intervals, which also will be the subject of future research.
@@
Another relevant application is the evaluation of HCD with individual (non-aggregated) values reported (e.g., for individual wells in one historical MNT-experiment rather than the complete plate), where extra-Binomial variability is also possible within one experiment (e.g. between the wells of the same plate). Here, GLMMs (both frequentist and Bayesian) as well as hierarchical models can be very helpful. Jeske and Young 2013 provided a bootstrap based procedure for the calculation of prediction intervals based on one-way GLMM, but to the authors knowledge, a general solution for the calculation of prediction intervals drawn from GLMM is not yet available. Hence, also this topic remains as an open issue that needs further investigation.
@@ 8 Conclusions
  • Prediction intervals computed based on Bayesian GLMM have coverage probabilities which are relatively close to the nominal level (in most cases). Because they can be easily adapted to handle more than one level of hierarchy, they are promising candidates for the application to non-aggregated HCD.
@@ 9 Acknowledgments
We have to thank Frank Schaarschmidt for fruitful discussions on prediction intervals and the impact of overdispersion on the inference drawn from dichotomous data. Furthermore, we have to thank both reviewers for their valuable comments which greatly helped to improve our manuscript.

### v4

@@ 5.2 Coverage probabilities for the simulation setting inspired by long term carcinogenicity studies
The heuristical methods (Fig. 4) behaved in a similar fashion than in the MNT-setting: With rising amount of historical control groups, the future observation is always covered (Fig. 4A). In the absence of overdispersion, the np-Chart (Fig. 4B) yields coverage probabilities close to the nominal level, but with a rising amount of overdispersion, its coverage probabilities decrease below 0.6. With a rising amount of historical control groups (at least 20), the coverage probabilities of the HCL computed by the mean $\pm$ 2 SD (Fig. 4C) approach the nominal level, if the underlying distribution is relatively symmetrical (low overdispersion and relatively high proportion). With increasing right-skewness of the underlying distribution (increasing overdispersion and decreasing proportion) and /or a low amount of historical control groups ($H<20$) the HCL computed by the mean $\pm$ 2 SD yield coverage probabilities below the nominal level that can decrease to approximately 85 %.
@@ 6 Application of control limits
Table 3: Historical control limits for the mortality of male B6C3F1-mice obtained in negative controls of long term carcinogenicity studies at the National Toxicology Program.
@@ 7 Discussion
The validation of a current control group based on prediction intervals is recommended by several authors (Menssen 2023, Kluxen et al. 2021) and since shortly, seems to be preferred by the European Food Safety Agency (EFSA 2024) over other (heuristical) methods. Our simulations have shown that heuristical methods do not control the statistical error to a satisfactory level and hence, can not be recommended for the practical validation of a current control group. Contrary to this, we could show, that three of the four proposed prediction intervals satisfactorily control the statistical error, if calculated based on a model that match the statistical properties of the HCD.
7.1 Bayesian prediction intervals
@@
Contrary to the Bayesian hierarchical models, the Bayesian GLMM did not have convergence problems in the LTC-setting. But, the applied GLMM mainly yield prediction intervals with coverage probabilities above the nominal level (except for a higher amount of overdispersion). This is, because such models a priori assume between-study variability to be present, and hence provide estimates $>0$ for the respective parameters, even if between-study variability is absent in the the data-generating process.
7.2 Bootstrap calibrated prediction intervals
@@
The coverage probability of the calibrated intervals is mainly effected by the amount of available historical information, defined by the numbers of experimental units within each control group $n_{h}$ and number of historical studies $H$. In most of the cases the amount of overdispersion or the magnitude of the binomial proportion has minor impact on the coverage probability, as long as one accepts that the lower border can become zero for low proportions and / or high overdispersion such that the desired 95 % prediction interval practically results in a 97.5 % upper prediction limit.
For a high cluster size $n_{h}$ (as in the MNT-setting) both calibrated prediction intervals yield coverage probabilities satisfactorily close to the nominal level, even for only five historical control groups. Contrary, for smaller cluster sizes (as in the LTC-setting) the number of historical control groups has to be slightly higher (say 10) to enable coverage probabilities close enough to the nominal level. This difference can be explained by the fact, that with increasing numbers of experimental units within each control group ($n_{h}\to\infty$) the available information on the success rates $\pi_{h}$ rises. Consequently also their estimates $\hat{\pi}_{h}$ become more precise. With other words, the precision of $\hat{\pi}_{h}$ is practically high enough to properly estimate the magnitude of the between-study overdispersion if $n_{h}=18000$ even for $H=5$. Contrary, if $n_{h}=50$ (or even lower), the information needed for the estimation of $\hat{\pi}_{h}$ carried by the data is heavily limited. Especially if the true (but unknown) success rate $\pi$ also gets low, the data contains many zeros. This leads to two implications:
First, the estimates $\hat{\pi}_{h}$ become heavily imprecise. Second, the estimates for the quasi-binomial dispersion parameter $\phi$ as well as for the intra-class correlation $\rho$ are known to be negatively biased (McCullagh and Nelder 1989, Saha and Paul 2005). Consequently, it can happen, that the data practically appears to be underdispersed, even if it derives from a data generating process that is inherent of between-study variation (especially, if most of the $\hat{\pi}_{h}$ are estimated to be zero). Since underdispersion is biologically implausible (as explained in the last paragraph of section 2), it was necessary to restrict the estimate for the quasi binomial dispersion parameter to $max(\hat{\pi},1.001)$. Similarly, the estimate for the beta binomial intra-class correlation was restricted to be $max(\hat{\rho},0.00001)$. Note that both restricted estimates have comparable magnitudes of relative bias, but the estimate for the intra-class correlation has a higher coefficient of variation (especially if the number of available control groups $H$ is low, see section 6 of the supplementary material). These higher variability might explain, that the beta-binomial prediction interval tends to be slightly liberal for a low number of historical control groups ($H=5$), whereas the quasi-binomial prediction interval behaves in a conservative fashion for this scenario.
7.3 Practical handling of overdispersed data
The magnitude of overdispersion reflects the between-study variation in toxicological HCD. However, systematic between-study variation might be induced by a mixture of both, controllable causes such as different handling of experimental units by different lab personnel, as well as uncontrollable causes such as the genetic variation inherent in the population of experimental units. This means that in practice, overdispersed HCD should be handled with care in a way that the magnitude of overdispersion reflects mainly the biological variability inherent of the biological process under observation, rather than the variability that is induced by controllable causes such as different handling of experimental units by different lab personnel or the change of the supplier of the model organism (Dertinger et al. 2023). However, the acceptable magnitude of between-study overdispersion depends on the biology of the specific model organism used for a specific type of study. Consequently, a general advise on its acceptable magnitude can not be given here, but rather has to be discussed separately for each type of study.
7.4 Perspectives
@@
The bootstrap calibration procedure presented above was already applied to enable the calculation of prediction intervals drawn from linear random effects models (Menssen and Schaarschmdit 2022) or in order to predict overdispersed count data (Menssen et al. 2024). Its further application to ovderdispersed binomial data proved, that our algorithm is a general and flexible tool for the calculation of prediction intervals. Hence, for the future it is planned to exploit its potential to enable the calculation of prediction intervals for a broader range of models / distributions. Furthermore, it is possible to extend the algorithm to different types of simultaneous prediction, which will be the subject of future research.
@@
A highly relevant application not yet covered by this manuscript is the evaluation of HCD on the base of individual (non-aggregated) values (e.g., for the number of micronuclei for each of the six individual wells in one historical MNT-experiment rather than for the sum of micronuclei per plate) where extra-binomial variability is also possible within one experiment (e.g. between the wells of the same plate). For this type of data, generalized linear mixed models need to be applied. Jeske and Young 2013 provided a bootstrap based procedure for the calculation of prediction intervals based on one-way GLMMs, but to the authors knowledge, a general solution for the calculation of prediction intervals drawn from GLMMs is not yet available. In the current manuscript we showed how to derive prediction intervals from a Bayesian GLMM with one level of hierarchy. However, this approach can easily be extended to more complex designs, but these extensions will need further investigations.
@@ 8 Conclusions
  • Prediction intervals computed based on Bayesian GLMM have coverage probabilities which are relatively close to the nominal level (in most cases). Because they can be easily adapted to handle more than one level of hierarchy, they are promising candidates for the application to HCD that follows a more complex design.
@@ 9 Acknowledgments
We have to thank Frank Schaarschmidt for fruitful discussions on prediction intervals and the impact of overdispersion on the inference drawn from dichotomous data. Furthermore, we have to thank all persons involved in the reviewprocess for their efforts and their valuable comments in three rounds of comprehensive review.